OpenStudy (anonymous):

Eliminate the parameter. x = t^2 + 3, y = t^2 - 2

OpenStudy (anonymous):

@e.mccormick @satellite73

OpenStudy (anonymous):

@agreene @Best_Mathematician @Christos @Danny_Boy @enya.gold @flutterflies

OpenStudy (e.mccormick):

Solve for t, set themm both = to each other. Then you have it in x and y. That is probably what they mean, because t is generally the parameter.

OpenStudy (anonymous):

the answer options are y = x + 5, x ≥ 3 y = x2 - 5, x ≥ 3 y = x - 5, x ≥ 3 y = x2 + 5, x ≥ 3

OpenStudy (e.mccormick):

OK. So they are having you unparamiterize it that way. This in Linear?

OpenStudy (anonymous):

yes i think so

OpenStudy (e.mccormick):

\[x = t^2 + 3\\ y = t^2 - 2\\ \therefore \\ \left[\matrix{ 1 & 0 & | t^2 + 3\\ 0 & 1 & | t^2 - 2 }\right]\Rightarrow \left[\matrix{ 1 & -t^2 & |\quad 3\\ -t^2& 1 & | - 2 }\right]\]

OpenStudy (e.mccormick):

That would be one way of working towards it in Linear.... Which lets you see the relationships it started from. What you would be looking for as an answer is the solution space where that holds true.

OpenStudy (e.mccormick):

Hmmm... Well, might not be quite the best matix to use to represent this. However, it does let me come up with an idea.

OpenStudy (e.mccormick):

\(x-3=t^2\) And: \(y+2=t^2\), but \(t^2=t^2\), \(\therefore \, x-3 = y+2\) Now to see how those relate to: y = x + 5, x ≥ 3 y = x2 - 5, x ≥ 3 y = x - 5, x ≥ 3 y = x2 + 5, x ≥ 3

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

Well, the \(x−3=y+2\Rightarrow x−1=y\) Ahh. I think I am losing answers because I did not square it. because I do not see that anywhere in the list.

OpenStudy (e.mccormick):

I did a little search to see why I did not remember this in the linear I know. It is trig! Not linear. You in a trig class?

OpenStudy (e.mccormick):

Or it can be...

OpenStudy (anonymous):

im in pre calc we are doing trig right now

OpenStudy (e.mccormick):

Kk. So this is most likely for trig. Now I see why I did not recognize the form right off. I took Trig and Alg, but skipped pre-calc. Probably something they did in there. I got some really ugly answers and some simple ones, but I did not get one that matched up with the possible choices, so I know I am doing something wrong. @Mertsj @jim_thompson5910 Anyone care to dig out of this hole I found? I see the point of this, but using what I know does not give me any of the 4 answers listed, so I am obviously off somewhere.

jimthompson5910 (jim_thompson5910):

why not isolate t^2 instead of isolate t?

jimthompson5910 (jim_thompson5910):

that's much easier in my opinion

OpenStudy (mertsj):

t^2=x-3 t^2=y+2 y+2=x-3 y=x-5

OpenStudy (e.mccormick):

I tried that, but it did not get me where I wanted, so I must have made a mistake somewhere.

jimthompson5910 (jim_thompson5910):

x = t^2 + 3, y = t^2 - 2 --------------- x = t^2 + 3 x-3 = t^2 t^2 = x-3 --------------- y = t^2 - 2 y = (x-3) - 2 y = x - 3 - 2 y = x - 5

OpenStudy (e.mccormick):

AH HA! I made a simple algebra error. LOL! I got \(x-1\) and my brain got stuck there!

OpenStudy (anonymous):

so it would be the third option?

OpenStudy (e.mccormick):

Thanks guys! @mgarcia634 OK... so if you do not make the stupid mistake I made... you get what Jim did. it is exactly the same thing Mertsj started and I did earlier when I did: \[x-3 = t^2\\ y+2 = t^2\\ x-3 = y+2 \] Then I went and added 2 rather than subtract it, which is where I confused the heck out of myself, and thereby confused you.

OpenStudy (e.mccormick):

Yep! Third it is.

OpenStudy (anonymous):

thanks guys!

OpenStudy (e.mccormick):

Really sorry about the confusion.

OpenStudy (anonymous):

it's fine